OpenStudy (anonymous):

The height of a helicopter above the ground is given by h = 2.85t3, where h is in meters and t is in seconds. At t = 1.65 s, the helicopter releases a small mailbag. How long after its release does the mailbag reach the ground?

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

i mean i am looking for the answer

OpenStudy (anonymous):

later

OpenStudy (anonymous):

\[a=10m s ^{-2}\]\[u=0 \] s=2.85m \[v^2-u^2=2as\] \[v^2= 20m/s^2 * 2.85\] \[v^2= 57m^2 s ^{-2}\] \[\sqrt{57} = 7.55\] \[a=v/t [u=o]\] \[10m/s^2 = 7.55m/s \div t\]\[t= 5.77/10 = 0.57s\] the time it reaches the ground from releasing from the sky = 0.57s after the heli take off is 1.65+0.57s =2.22s